A car of mass M = 1400 kg traveling at 40.0 km/hour enters a banked turn covered with ice. The road is banked at an angle ∘, and there is no friction between the road and the car's tires. Use g = 9.80 m/s2 throughout this problem.

What is the radius r of the turn if θ = 20.0 ∘ (assuming the car continues in uniform circular motion around the turn)? Express Answer in meters.
Answer #1

FNsinθ = ma​​​​​c

Since a​​​​​​c = v​​​​​​2/r:

FNsinθ = mv​​​​​2/r.....(1)

Also, we know that,

FNcosθ = mg........(2)

If we divide the equation (1) by (2)

((FNsinθ) / (FNcosθ)) = (mv​​​​​2/r) / (mg)

Which gives us:

tanθ = (mv​​​​​2/r) / (mg)

simplified to:

tanθ = v​​​​​​2 / (rg)

Now,

40km/h = 40000m/h or,

= 11.11 m/s

Next use the formula we solved for:

tanθ = v​​​​​​2 / (rg)
tan(20) = 11.112/9.81*r
0.364 = (12.582/r)

r = 34.566 m